A father is 26years older than his daughter. In 4years he will be 2 more than 5x her age. What are their current ages?

Answer by Maths68(1474)

Let
Present age of father = f
Present age of daughter = d

A father is 26years older than his daughter.
f=d+26...........(1)

In 4 years
Father will be = f+4
Daughter will be = d+4
Then
father will be 2 more than 5 times of her age.
f+4=5(d+4)+2
f+4=5d+20+2
f+4=5d+22
f=5d+22-4
f=5d+18.................(2)
Put the value of f from (1) to (2)
d+26=5d+18
d-5d=18-26
-4d=-8
Divide by -4 both sides
-4d/-4=-8/-4
d=2
Put the value of d in (1)
f=d+26
f=2+26
f=28

Present age of father = f = 28 years old
Present age of daughter = d = 2 years old